OpenStudy (daniellelovee):

Suppose the dial on the spinner is spun 2 times in a row. Create a probability histogram for the variable X.

OpenStudy (daniellelovee):

@Directrix

OpenStudy (daniellelovee):

page 151

OpenStudy (daniellelovee):

@dan815

OpenStudy (anonymous):

question on page 151 about the spinner?

OpenStudy (daniellelovee):

yup but the question I posted

OpenStudy (daniellelovee):

X is the number of times the dial lands on section A.

OpenStudy (anonymous):

probability histgram for the number of times it lands on A? is that the question?

OpenStudy (daniellelovee):

yup

OpenStudy (anonymous):

so there are 3 possible numbers for X it could land 0 times, 1 time or 2 times

OpenStudy (anonymous):



OpenStudy (anonymous):

now we need the probability of each of these do you know what the probability you get something other than A both times?

OpenStudy (anonymous):

if you do not, that is fine, i will tell you just asking

OpenStudy (daniellelovee):

is it 4/9?

OpenStudy (anonymous):

yes it is !

OpenStudy (daniellelovee):

ok so far so good then lol

OpenStudy (daniellelovee):

ok and for 2 is it 2/9?

OpenStudy (anonymous):

how about the probabilty it lands on A two times (i skipped one i know)

OpenStudy (daniellelovee):

sorry I meant for 1

OpenStudy (daniellelovee):

ok hold on lol for 0 is 4/9 for 1 is 2/9 and for 2 is 1/9?

OpenStudy (anonymous):

not quite the first and last are right,that is why i skipped 1

OpenStudy (anonymous):

you computed \[\frac{2}{9}\]by \[\frac{1}{3}\times \frac{2}{9}\] right?

OpenStudy (daniellelovee):

yeah

OpenStudy (anonymous):

but that would be the probability it lands A on the first spin, not A on the second we have another possibility

OpenStudy (anonymous):

namely not A, then A of course they have the same probability, but they are two disjoint events, so you have to add them up

OpenStudy (daniellelovee):

so then would it be the same as 0 based on what you are saying? it would be 4/9 again?

OpenStudy (anonymous):

also notice that it has to land on A 0, 1 or 2 times but the probabilities need to add up to 1, and they don't the way you wrote it

OpenStudy (anonymous):

yeah \[\frac{4}{9}\]

OpenStudy (daniellelovee):

ok so the 4/9, 4/9, and 1/9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and they do add to 1 right?

OpenStudy (daniellelovee):

yes

OpenStudy (daniellelovee):

thank you so much for your help and can I ask one more?

OpenStudy (anonymous):

then it looks good i will let you draw it yourself, make sure that \(\frac{1}{9}\) is only one fourth as high as \(\frac[4}{9}\) sure ask away

OpenStudy (anonymous):

but let me ask a question too where did you get this text, and what grade is it for?

OpenStudy (daniellelovee):

ok and for that is ok I needed the table and from there I can make the graph :) and for the secon one I dont get is from page 150 number 8

OpenStudy (anonymous):

it will take a minute to find it hold on a sec

OpenStudy (daniellelovee):

ok is easier if you use ctrl+F and type 2 trials

OpenStudy (anonymous):

i got it, but 8 means 7,8,9,10 right? four bar graphs or histogram or whatever

OpenStudy (daniellelovee):

yeah but I just need 8 :)

OpenStudy (daniellelovee):

we are doing even numbers and some odd numbers

OpenStudy (daniellelovee):

and im only missing those 2 from that lesson :)

OpenStudy (anonymous):

you can't really do just one, but no matter in fact it is better if we do them all by sheer reason rather than actually making the histogram, although we can do that if you like

OpenStudy (daniellelovee):

I meant as in questions numbers not from the graph

OpenStudy (anonymous):

maybe it is not clear, but maybe it is, that the higher the probability of success, the more the histogram is skewed to the right

OpenStudy (anonymous):

is it "match the binomial probability to the histogram" question at the bottom of the page?

OpenStudy (daniellelovee):

Match the probability histogram with the description of the binomial distribution. Assume that there are 2 trials.

OpenStudy (anonymous):

right so we can't really do 8, you have to match them doing 8 doesn't really make sense since 8 is a graph

OpenStudy (anonymous):

lets to them all without doing them the higher the probability of success, the more likely you will have successes, the more the histogram will be skewed to the right (lower in the 0, higher in the 2)

OpenStudy (daniellelovee):

I dont understand that sorry :(

OpenStudy (daniellelovee):

I dont understand that sorry :(

OpenStudy (anonymous):

if i get a success with probability \(.8\) then the probability i get two successes is \(.8\times .8=.64\) but the probability i get no successes is \(.2\times .4=.04\) a much smaller number

OpenStudy (anonymous):

typo there, i meant \(.2\times .2=.04\)

OpenStudy (anonymous):

so the graph will be higher on the right, lower on the left

OpenStudy (daniellelovee):

I meant the graph from number 8 from the 8th question

OpenStudy (anonymous):

is that clear more or less

OpenStudy (daniellelovee):

but how did you get probability 0.8?

OpenStudy (anonymous):

ok lets go slow you have to match them up it will be very very hard to look at the histogram and decide which of the four binomials it is

OpenStudy (anonymous):

i made up the .8 as an example

OpenStudy (daniellelovee):

im soo confused with this sorry

OpenStudy (daniellelovee):

oh ok got it

OpenStudy (anonymous):

i think you might be confused as to what the question is asking you cannot really do just 8 you have to match them

OpenStudy (anonymous):

you have A, B, C, D and 7, 8, 9, 10 and you are supposed to say A goes with 10 for example and B goes with ...

OpenStudy (daniellelovee):

ohhh ok match them to the letter questions above?

OpenStudy (daniellelovee):

ok got it now

OpenStudy (anonymous):

RIGHHHHHT

OpenStudy (anonymous):

ok so now that we know what we have to do, lets see if we can do it

OpenStudy (anonymous):

you don't really want to make four histograms for four different binomial probabilities just to see what you get do you? too much work

OpenStudy (daniellelovee):

no lol too lazy

OpenStudy (anonymous):

so lets to it the simple way and see if we can understand it as well

OpenStudy (daniellelovee):

alright

OpenStudy (anonymous):

the higher the probability is, the more likely you get a success (that is really just saying larger probability is larger, no real content there) which one of A, B, C, D has the highest probability of success?

OpenStudy (daniellelovee):

A

OpenStudy (anonymous):

and which of those 4 graphs has the highest number for two successes and the lowest number for 0

OpenStudy (daniellelovee):

but I think it would be 0.3 based on what you are saying because when there 2 trials graph 8 is the 3rd one fron big to small probability

OpenStudy (anonymous):

yes i think so too

OpenStudy (anonymous):

since you seem to be focused only on 8 i would say A goes to 10 B to 7 etc you agree?

OpenStudy (daniellelovee):

alright would you mind if I got a second opinion on this because it seems too easy 0.0

OpenStudy (anonymous):

no not at all i go to two doctors we can actually check that \(.3\) gives the graph of 8 if you like

OpenStudy (anonymous):

zero success is \(.7\times .7=.49\) and two successes is \(.3\times .3=.09\)

OpenStudy (anonymous):

which looks pretty much like that graph right?

OpenStudy (daniellelovee):

yeah

OpenStudy (anonymous):

one success is \(2\times .3\times .7=.42\)also looksgood

OpenStudy (daniellelovee):

yup

OpenStudy (anonymous):

ok so i think we nailed that one let me ask again, what book, what grade?

OpenStudy (daniellelovee):

and thank you so much for you help :)

OpenStudy (anonymous):

yw because i like it, seems easy enough to understand

OpenStudy (anonymous):

you done with these?